The internal energy change for reaction is -885j k mol-1 at 300 k.calculate its enthalpy change.

Grade:11

1 Answers

Mayank
4 years ago
using the formulae H=U + ngRT  where H=entahlpy change U=internal energy change ng=one mole of the reactant converted into moles of product  so assuming in rxn ng= 1  we get
H=U +RT  putting in the values we get
H=-885+8.314*300= 1609.2jk mol-1                                // note R=8.314